OpenStudy (anonymous):

can anyone please double check my work? the math problem is ..write an equation of the line containing the given point and perpendicular to the given line..... (0,9);2x+7y...I got the answer of -2/9x+9

OpenStudy (amistre64):

a perp.line has the propety that the slope of the lines, when multiplied together will be (-1)

OpenStudy (amistre64):

y=-2/7x +.... perp line: y = 7/2x + b (x=0,y=9) 9 = 7/2(0) + b b=9 perp line: y = 7/2x + 9 {9 = 18/2} -7x + 2y = 18 ; some texts dont like the (-) in front so multiply everything by (-1) 7x - 2y = -18

OpenStudy (amistre64):

(-2/7)(7/2) = -1

OpenStudy (amistre64):

think of the slope as: -(first)/(last) which in this case is: -(2)/(7); now we need to multiply this by some number to get (-1) -(2/7)n = -1 -2n = -7 n = -7/-2 ; since 2negatives give a positive we write: n = 7/2 for the slope of the perp.line...